OpenStudy (anonymous):

In Homework Handout 1.0, section 1.4 (#4) how do you find the nth root? I know to square an integer you type **2, to find the square root you type **.5, but how about the nth? (n = 4 )

OpenStudy (e.mccormick):

They are talking about using one of the classical methods. For using exponents, look at .5 another way. \(.5=\frac{1}{2}\) and in roots, \(\Large\sqrt[a]{x}=x^{\frac{1}{a}}\)

OpenStudy (e.mccormick):

Classical methods: http://en.wikipedia.org/wiki/Newton%27s_method http://www.youtube.com/watch?v=ZMFMdggWTaw That is another way of finding roots where you have to see if you are close enough to call it good. That is the whole epsilon thing in the lectures. If a human does it, you decide if it is close enough by instinct or by math. For a computer, you need to set up a conditional test and do it by math.

OpenStudy (anonymous):

wonderful, thank you for the fast clarification and informative videos!!